# Characterising the Creative Process in Humans and Large Language Models

### v1

@@ Introduction
Much recent work has benchmarked and quantified the generative creative aptitudes of large language models (LLMs) (??????). LLMs often perform as well as the average human on creative thinking tasks such as the Alternate Uses Task (AUT) (?????). However, these works largely analysed creativity from a Product perspective (?), assessing how original and useful model responses are to determine “what makes them creative (or not)”. An equally important component of creativity, less studied in the field of Artificial Creativity, is the Process perspective (?), addressing the question of “how creativity arises”. This paper aims to fill this gap and characterise human and LLM creativity by looking at the creative process (?), particularly the way humans and LLMs explore semantic spaces while generating creative ideas.
@@
In the next sections, we first introduce method and investigate its reliability and validity. We then apply it to characterise human and LLM flexibility on the AUT and VFT. We find that LLMs as a population match the variability in human response sequences on AUTs, but unlike humans, their relationship to creativity differs. We also discuss how to use these insights to use LLMs as artificial participants or co-creators.
@@
Human data were collected from (anonymized) undergraduate participants using a within-subjects design. For the AUT, participants listed as many creative uses for “brick” and “paperclip” as possible in a fixed time of 10 minutes. For the VFT, participants named as many animals as possible in a fixed time of 2 minutes. Participants not adhering to instructions were removed, resulting in a total of 220 participants. The responses, originally in Dutch were translated to English for analysis using the deep-translator Python package. Translations were manually inspected to correct for errors due to spelling mistakes.
@@ Results
We first test the reliability and validity of the jump signal. For reliability, we measured the test-retest correlation of the number of jumps for AUT brick and paperclip response sequences from 81 participants (who had ${>}{=}{18}$ responses in both). We found a positive Pearson correlation of ${r}{=}{0.42}$ (${p}{<}{0.001}$, ${CI}{=}{[0.22,0.58]}$), which is high considering the test-retest and alternate-form reliability of AUT product creativity seldom exceeds ${r}{=}{0.5}$.
@@ Discussion
Our work has limitations. First, we used the same embedding model (with different distance metrics) for response categorisation and semantic similarities. Second, each response was categorised into a single category—however, a response such as “throw a brick to produce sound” should include concepts of both “throw” and “produce sound”. Multiclass classification based on predefined categories could tackle this issue. Third, we only scored response originality ignoring utility, which could inflate creativity comparisons as inappropriate responses (e.g., “using a brick as a ribbon”) have high originality scores but are not considered creative. Lastly, the AUT being a popular creativity task (especially AUT brick333checked with WIMBD tool (?)) could be a source of LLM data contamination (?).

### v2

@@ Introduction
Much recent work has benchmarked and quantified the generative creative aptitudes of large language models (LLMs) (??????). LLMs often perform as well as the average human on creative thinking tasks such as the Alternate Uses Task (AUT) (?????). However, these works largely analysed creativity from a Product perspective (?), assessing how original and useful model responses are to determine “what makes them creative (or not)”. An equally important component of creativity, less studied in the field of Artificial Creativity, is the Process perspective (?), addressing the question of “how creativity arises”. This paper aims to fill this gap and characterise human and LLM creativity by looking at the creative process (?), comparing the way humans and LLMs explore semantic spaces while generating sequences of creative ideas. This comparison makes sense given that the tasks involve creative response-sequence generation, in line with LLM objective of next-word prediction.
@@
In the next sections, we first introduce our method and investigate its reliability and validity. We then apply it to characterise human and LLM flexibility on the AUT and VFT. We find that LLMs as a population match the variability in human response sequences on AUTs, but unlike humans, their relationship to creativity differs. We also discuss how to use these insights to use LLMs as artificial participants or co-creators.
@@
Human data were collected from undergraduate participants at University of Amsterdam using a within-subjects design. For AUT, participants listed as many creative uses for “brick” and “paperclip” as possible in a fixed time of 10 minutes. For VFT, participants named as many animals as possible in a fixed time of 2 minutes. Participants not adhering to instructions were removed, resulting in a total of 220 participants. The responses, originally in Dutch were translated to English for analysis using the deep-translator Python package. Translations were manually inspected to correct for errors due to spelling mistakes.
@@ Results
We first test the reliability and validity of the jump signal. For reliability, we measured the test-retest correlation of the number of jumps for AUT brick and paperclip response sequences from 81 participants (who had ${>}{=}{18}$ responses in both). We found a positive Pearson correlation of ${r}{=}{0.42}$ (${p}{<}{0.001}$, ${CI}{=}{[0.22,0.58]}$), which is high considering the test-retest and alternate-form reliability of AUT product creativity seldom exceeds ${r}{=}{0.5}$ (?).
@@ Discussion
Our work has limitations to be addressed with further research. First, we used the same embedding model (with different distance metrics) for response categorisation and semantic similarities. Second, each response was categorised into a single category—however, a response such as “throw a brick to produce sound” should include concepts of both “throw” and “produce sound”. Multiclass classification based on predefined categories could tackle this issue. Third, we used a very basic prompt for LLMs in order to match human instructions. Future work can experiment with more advanced prompt engineering. Fourth, we only scored response originality ignoring utility, which could inflate creativity comparisons as inappropriate responses (e.g., “using a brick as a ribbon”) have high originality scores but are not considered creative. Lastly, the AUT being a popular creativity task (especially AUT brick333checked with WIMBD tool (?)) could be a source of LLM data contamination (?).
@@
Acknowledgements
Authors thank Fleur Doorman and Emma Schreurs for their Masters theses work focussing on aspects of this project. Nath thanks Vishaal Udandarao, Andrew Webb, Sahiti Chebolu for useful suggestions and code review. The data collection and previous work on this project was funded by the Amsterdam Brain & Cognition (ABC) Talent Grant (University of Amsterdam) 2016–2018 awarded to Stevenson and the Jacobs Foundation Fellowship 2019–2022 awarded to Stevenson (2018 1288 12).